OpenStudy (anonymous):

A person jumps from the roof of a house 3.1-m high. When he strikes the ground below, he bends his knees so that his torso decelerates over an approximate distance of 0.70 m. If the mass of his torso (excluding legs) is 42kg . Part A Find his velocity just before his feet strike the ground. Express your answer using two significant figures. Enter positive value if the velocity is upward and negative value if the velocity is downward.

OpenStudy (anonymous):

Part B Find the average force exerted on his torso by his legs during deceleration. Express your answer using two significant figures. Enter positive value if the force is upward and negative value if the force is downward.

OpenStudy (anonymous):

i got an answer of -6.9 m/s but it wrong

OpenStudy (anonymous):

that was for part a and part b i got 33.6 m/s2

OpenStudy (fifciol):

to get the answer for A simply use conservation of mechanical energy mgh=mv^2/2 v=sqrt(2gh) pointing down

OpenStudy (anonymous):

ahh thank you

OpenStudy (anonymous):

i got 7.79, how would i do part b

OpenStudy (anonymous):

i got an answer of 1820.523 N i dont know if i am correct

OpenStudy (xishem):

One of the one-dimensional motion equations is relevant here:\[v^2=v_0^2+2ax\]Doing some algebra, plugging in numbers, and using Newton's second law, you'll find the answer you got: +1821N. Remember, though, that your reported velocity should be negative, since it's in the "down" direction. Also, report using the correct number of significant figures.

OpenStudy (anonymous):

i have a question would i use the x of 0.70 or 3.1 m

OpenStudy (xishem):

The equation should really be written like this:\[v_f^2=v_i^2+2a(\Delta x)=v_i^2+2a(x_f-x_i)\]The v_f and the x_f are corresponding values. Whatever the velocity you used for v_f should have a corresponding position, x_f. The same goes for v_i and x_i. Since you're looking at the time between him hitting the ground and completely stopping moving, you are looking at 0.70m (x_i) and 0m (x_f) as well as -7.79m/s (v_i) and 0m/s (v_f). You just need to be sure to use positions and velocities that correspond with each other if that makes sense... Maybe my explanation isn't the best.

OpenStudy (anonymous):

no i do get it, thank you so much =)

OpenStudy (anonymous):

and if its in 2 sig figs it would be 1.8 x 10^3

OpenStudy (xishem):

That's right (:.

OpenStudy (anonymous):

it saying my answer is wrong

OpenStudy (anonymous):

i dont know what i did wrong

OpenStudy (xishem):

Did you include units?

OpenStudy (anonymous):

yes i put N

OpenStudy (xishem):

Was your velocity value correct?

OpenStudy (anonymous):

yes i used 7.8

OpenStudy (anonymous):

well -7.8

OpenStudy (xishem):

I'm afraid I don't know. The best I can say is to try reporting it as 1800N, but otherwise either we're wrong or the program is just not accepting that answer for whatever reason.

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Physics?

OpenStudy (anonymous):

yes

OpenStudy (luigi0210):

I have never taken physics in my life: @agent0smith

OpenStudy (anonymous):

oki thats fine thank u

OpenStudy (agent0smith):

haha i solved this problem for my AP physics class a few days ago.

OpenStudy (anonymous):

i tried so many times to figure this out but i got it wrong

OpenStudy (agent0smith):

Oh in my book it was 3.9m tall... are you sure about the 3.1m? Its the Giancoli Physics book, 6th Ed.

OpenStudy (agent0smith):

All other numbers are the same, except 3.9 m not 3.1 m.

OpenStudy (anonymous):

i have the 7th edition

OpenStudy (anonymous):

and this is homework we have to submit online

OpenStudy (agent0smith):

Oh, well... then idk, i don't see why 1.8x10^3 N is wrong :/

OpenStudy (anonymous):

hmm how did u do your problem out?

OpenStudy (agent0smith):

Same way you and @Xishem did, using the same equations and F=ma.

OpenStudy (anonymous):

idk what distance to use?

OpenStudy (anonymous):

would it be 0.70 or 3.1 for part b

OpenStudy (agent0smith):

for part b, you know the velocity, and that his torso decelerates over 0.7m. Use it to find accel.

OpenStudy (anonymous):

i did that then i plugged that into f=ma and got the force

OpenStudy (agent0smith):

Yep, idk why 1800N is wrong. I got that.

OpenStudy (agent0smith):

You could always try... 2.3x10^3 N, i thinkk that was the answer i got from my book's question (it's an even question and the answers only show odd questions so i can't actually check it). But i can't imagine why they'd update the velocity answer but not the force answer

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

can you help me out on another problem?

OpenStudy (anonymous):

i can post it in another section

OpenStudy (agent0smith):

ok

OpenStudy (agent0smith):

@Xishem we forgot that the NET force is what decelerates the torso, but this isn't the average force exerted ON his torso BY his legs during deceleration.